# Bayesian RG Flow in Neural Network Field Theories

### v2

@@ 2.2 Bayesian Inference for Neural Networks
Once the NN architecture, including its initialization distribution $\pi_{0}$, has been specified one can begin learning. The training data is presented as a set of matched pairs, $D\equiv\{x_{i},\phi_{i}\}_{i\in\mathcal{I}}$ where $x_{i}\in I$ and $\phi_{i}\in O$. Typically, training a NN is understood as a problem in convex optimization whereupon the network is asked to tune its parameters in order to minimize a prescribed loss function $L(D\mid\theta)$. To obtain robust parameter estimates, one performs training on an ensemble of networks with values initialized from the distribution $\pi_{0}$. The resulting initialized models are subsequently trained using a method of choice, such as gradient descent. The result of training is to acquire a set of parameter estimates, one for each network in the ensemble, which we denote by $\{\theta_{n}(D)\}_{n=1}^{K}$. One may interpret $\{\theta_{n}(D)\}_{n=1}^{K}$ as draws from some new parameter distribution identifiable with a Bayesian posterior, $\pi_{D}$. Thus, in this sense, network training can be interepted according to the same arrow (16) which characterizes Bayesian inference. These estimates are subsequently pooled to arrive at an optimal $\theta_{*}(D)$ resulting in the trained network $\phi_{\theta_{*}(D)}$.
@@
The initialization distribution $\pi_{0}$ now plays the role of a Bayesian prior. Given the data $D=\{(x_{i},\phi_{i})\}_{i\in\mathcal{I}}$ one can update the prior using Bayes’ law to obtain the distribution $\pi_{D}(\theta)$. From this perspective, rather than having an ensemble of networks we have promoted the parameters of network to the status of random variables. The gradient descent estimates $\{\theta_{n}(D)\}_{n=1}^{K}$ can now be regarded simply as draws from the posterior distribution, $\pi_{D}(\theta)$, and the optimal estimate is obtained by taking the maximum a posteriori estimate
@@ 2.3 Summary
Data generating distribution, $p_{*}$ Target function, $\phi_{*}$
@@
Sample, $D=\{y_{i}\}_{i\in\mathcal{I}}$ Training data, $D=\{x_{i},\phi_{i}\}_{i\in\mathcal{I}}$
@@ 3.1 Neural Network-Field Theory Correspondence
is the sampled $n$-point correlator for the distribution $p$; with an accuracy roughly proportional to the size of $T$.
@@
$p(\phi)=p_{0}(\phi)+\sum_{n=3}^{\infty}\frac{(-1)^{n}}{n!}\int_{I^{\times n}}\prod_{j=1}^{n}d^{d}x_{i}\;\;G^{(n)}(x_{1},...,x_{n})\frac{\delta^{n}p_{0}(\phi)}{\delta\phi(x_{1})\cdots\delta\phi(x_{n})}.$ (32)
@@
The standard result of Neal that an infinitely wide NN with independent and identically distributed parameters can be regarded as a Gaussian Random Process [3] is automatically encoded in the NNFT correspondence. It follows from the central limit theorem (CLT) which guarantees that all of the connected correlation functions of such a network, $G^{(n)}_{c}$, are zero for $n>2$. Thus, following (34), we see that $p(\phi)=\frac{1}{Z[0]}e^{-S_{\text{free}}(\phi)}$ coincides with a free field theory in this case. The power of NNFT truly presents itself when considering perturbations away from Neal’s seminal result, which may now be interpreted through the lens of perturbations away from a free field theory [32]. From this perspective, a natural way to generate non-Gaussianities (i.e. interactions) in an NNFT is to break the assumptions which are required for the central limit theorem to hold. Two straightforward ways of doing so are to move away from the strict $N\rightarrow\infty$ limit, or to choose a distribution $\pi$ in which parameters are no longer independent. The couplings appearing in $S_{\text{int}}$ 34 will have orders of magnitude associated with the extent to which the CLT is violated. For example, in the case that interactions arise from the finite width of the network while retaining the i.i.d. assumption over parameters it can be shown that the connected correlation functions scale as
$G^{(n)}_{c}(x_{1},\cdots,x_{n})\propto\frac{1}{N^{n/2-1}},$ (36)
leading to a $N$-scaling in the couplings.
@@ 4.1 Bayesian RG for NNGPs
At initialization, we take our network parameters to be distributed as [^18] We have introduced notation in which $i_{\ell}=1,...,n_{\ell}$ is an index for the space $O_{\ell}$ in a chosen basis.
@@
That is, the parameters in each layer are identically distributed, and each layer is independent from every other layer. The final output of the network is given by
@@ 4.2.1 Implementation and network architecture
$f(x_{1},x_{2}):=(a-x_{1})^{2}+b(x_{2}-x_{1}^{2}),$ (84)
The input in this case is two-dimensional, $x\in I=\mathbb{R}^{2}$, and is sampled from the Gaussian distribution $X\sim\mathcal{N}(0,\mathbb{I}_{2})$, where $\mathbb{I}_{2}$ is the standard $2\times 2$ identity matrix. Training is performed such that the output $y\in\mathbb{R}$ minimizes the $L^{2}$ loss function,
@@
where $\phi_{\theta}(x)$ is the network function parameterized by $\theta$. Optimization is performed using ‘vanilla’ gradient descent [^24] By ‘vanilla’ gradient descent, we mean traditional gradient descent as opposed to stochastic gradient descent..
It is typical in the training of NNs to start with a collection of random draws from a seed distribution over parameters, performing the chosen optimization procedure for each parameter draw in parallel. For our purposes, as was discussed in Section 2, it is natural to interpret the seed distribution as a prior over networks, and to treat the ensemble of parallel networks as realizations from the Bayesian posterior at each stage of training. In this way, we can approximate the posterior in a numerically tractable form by parsing calculable statistics from the ensemble. For our sample problem, we take the prior distribution to be Gaussian. Following the notation of Section 2, we denote the ensemble of parameters by $\{\theta_{n}(D)\}_{n=1}^{k}$, where $D=\{x_{i},\phi_{i}\}_{i\in\mathcal{I}}$ is a training sample. In this case, the training is supervised, so each network input datum, $x_{i}$, is matched with the corresponding ground truth model realization, $\phi_{i}\equiv f(x_{i})$.

### v3

@@ 2.1 Introduction to Bayesian Inference
To clarify some of the abstract concepts introduced in this section, let us consider a standard implementation of Bayesian inference. Suppose we are interested in determining the probability of heads for a possibly unfair coin. To perform inference in this example we collect a sample of $N$ tosses of the coin, $D=\{y_{i}\}_{i=1}^{N}$, where each $y_{i}$ belongs to the sample space $S=\{H,T\}$ and $H$, $T$ indicates the outcome of ‘head’ and ‘tail’, respectively. The data generating distribution of the coin is Bernoulli with probability of heads given by some $\theta_{*}\in[0,1]$. In particular, $p_{*}(y)=\text{Bern}_{\theta_{*}}(y)$ where here
$\text{Bern}_{\theta}(y=H)=\theta,\qquad\text{Bern}_{\theta}(y=T)=1-\theta.$ (17)
The natural model class for this inference problem is the space of Bernoulli distributions
$M=\{\text{Bern}_{\theta}\;|\;\theta\in[0,1]\}\simeq[0,1].$ (18)
This model space is equivalent to the interval $[0,1]$, which corresponds to the possible probabilities which the coin could assign to a head.
To perform Bayesian inference we must choose a prior probability on the space of models. For simplicity, let us take the flat prior,
$\pi_{0}(\theta)=1,\qquad\theta\in[0,1].$ (19)
This ensures that the probability that $\theta$ is contained in any subinterval of the model space is simply equal to the size of the interval. Intuitively, the flat prior places equal probability to all of the possible values of the parameter $\theta$. The likelihood in this example is equal to the probability of observing the given sample under the assumption that the coin has a probability $\theta$ of landing on heads. Specifically,
$\mathcal{L}(D\mid\theta)=\prod_{i=1}^{N}\text{Bern}_{\theta}(y=y_{i})=\theta^{N_{H}}(1-\theta)^{N_{T}},$ (20)
where $N_{H}$ ($N_{T}$) is the number of heads (tails) observed in the sample. The joint probability density for the sample and the parameter $\theta$ is given by
$p_{D,\theta}(D,\theta)=\theta^{N_{H}}(1-\theta)^{N_{T}}.$ (21)
Integrating this quantity over $\theta$, and applying Bayes’ law we obtain the normalized posterior
$\pi_{D}(\theta)=\frac{(N_{H}+N_{T}+1)!}{N_{H}!N_{T}!}\theta^{N_{H}}(1-\theta)^{N_{T}}=\text{Beta}_{N_{H}+1,N_{T}+1}(\theta).$ (22)
This is a Beta distribution with parameters $\alpha=N_{H}+1$ and $\beta=N_{T}+1$.
Such a distribution has a mean equal to $\frac{N_{H}+1}{N+2}$, which asymptotes to the true underlying parameter for sufficiently large samples. For finite samples, (22) is essentially centered [^10] The reason the posterior mean is not equal to $N_{H}/N$ is due to the effect of the prior. A flat prior on this model space can be interpreted as a Beta distribution with parameters $\alpha=\beta=1$. This has the quantitative effect of adding two additional samples; one head and one tail. on the empirical probability estimate. In this way, the Bayesian posterior reproduces the maximum likelihood estimate, but also allows for a direct computation of the probability that the true underlying parameter belongs to any interval in the model space. This allows, among other things, for the formation of probabilistic confidence intervals. [^11] It’s worth noting that the variance of the posterior distribution is given by $\text{Var}(\theta\mid D)=\frac{(N_{H}+1)(N_{T}+1)}{(N+2)^{2}(N+3)}$ (23) which scales as $1/N$ for large $N$. This means that the width of the posterior will becomes sharper around the maximum likelihood estimate as the sample becomes larger.
@@ 2.2 Bayesian Inference for Neural Networks
Once the NN architecture, including its initialization distribution $\pi_{0}$, has been specified one can begin learning. The training data is presented as a set of matched pairs, $D\equiv\{x_{i},\phi^{*}_{i}\}_{i\in\mathcal{I}}$ where $x_{i}\in I$ and $\phi^{*}_{i}\in O$. Here, $\phi^{*}_{i}$ is the value of the target function $\phi^{*}$ which we would like our NN to approximate evaluated at $x_{i}$, i.e. $\phi^{*}_{i}=\phi^{*}(x_{i})$. Typically, training a NN is understood as a problem in convex optimization whereupon the network is asked to tune its parameters in order to minimize a prescribed loss function $L(D\mid\theta)$. To obtain robust parameter estimates, one performs training on an ensemble of networks with values initialized from the distribution $\pi_{0}$. The resulting initialized models are subsequently trained using a method of choice, such as gradient descent. The result of training is to acquire a set of parameter estimates, one for each network in the ensemble, which we denote by $\{\theta_{n}(D)\}_{n=1}^{K}$. One may interpret $\{\theta_{n}(D)\}_{n=1}^{K}$ as draws from some new parameter distribution identifiable with a Bayesian posterior, $\pi_{D}$. Thus, in this sense, network training can be interepted according to the same arrow (16) which characterizes Bayesian inference. These estimates are subsequently pooled to arrive at an optimal $\theta_{*}(D)$ resulting in the trained network $\phi_{\theta_{*}(D)}$.
@@
The initialization distribution $\pi_{0}$ now plays the role of a Bayesian prior. Given the data $D=\{(x_{i},\phi^{*}_{i})\}_{i\in\mathcal{I}}$ one can update the prior using Bayes’ law to obtain the distribution $\pi_{D}(\theta)$. From this perspective, rather than having an ensemble of networks we have promoted the parameters of network to the status of random variables. The gradient descent estimates $\{\theta_{n}(D)\}_{n=1}^{K}$ can now be regarded simply as draws from the posterior distribution, $\pi_{D}(\theta)$, and the optimal estimate is obtained by taking the maximum a posteriori estimate
@@ 2.3 Summary
Data generating distribution, $p_{*}$ Target function, $\phi^{*}$
@@
Sample, $D=\{y_{i}\}_{i\in\mathcal{I}}$ Training data, $D=\{x_{i},\phi^{*}_{i}\}_{i\in\mathcal{I}}$
@@ 3.1 Neural Network-Field Theory Correspondence
where the tilde indicates that this is now the sampled estimate of the $n$-point correlator for the distribution $p$; with an accuracy roughly proportional to the size of $T$.
@@
$p(\phi)=p_{0}(\phi)+\sum_{n=3}^{\infty}\frac{(-1)^{n}}{n!}\int_{I^{\times n}}\prod_{i=1}^{n}d^{d}x_{i}\;\;G^{(n)}(x_{1},...,x_{n})\frac{\delta^{n}p_{0}(\phi)}{\delta\phi(x_{1})\cdots\delta\phi(x_{n})}.$ (39)
@@
The standard result of Neal that an infinitely wide NN with independent and identically distributed parameters can be regarded as a Gaussian Random Process [3] is automatically encoded in the NNFT correspondence. It follows from the central limit theorem (CLT) which guarantees that all of the connected correlation functions of such a network, $G^{(n)}_{c}$, are zero for $n>2$. Thus, following (41), we see that $p(\phi)=\frac{1}{Z[0]}e^{-S_{\text{free}}(\phi)}$ coincides with a free field theory in this case. The power of NNFT truly presents itself when considering perturbations away from Neal’s seminal result, which may now be interpreted through the lens of perturbations away from a free field theory [32]. From this perspective, a natural way to generate non-Gaussianities (i.e. interactions) in an NNFT is to break the assumptions which are required for the central limit theorem to hold. Two straightforward ways of doing so are to move away from the strict $N\rightarrow\infty$ limit, or to choose a distribution $\pi$ in which parameters are no longer independent. The couplings appearing in $S_{\text{int}}$ (42) will have orders of magnitude associated with the extent to which the CLT is violated. For example, in the case that interactions arise from the finite width of the network while retaining the i.i.d. assumption over parameters it can be shown that the connected correlation functions scale as
$G^{(n)}_{c}(x_{1},\cdots,x_{n})\propto\frac{1}{N^{n/2-1}}.$ (43)
The couplings, $g^{(r)}(x_{1},...,x_{r})$, can, in turn, be expressed as a sum of all connected $r$-point Feynman diagrams which are constructed with $G_{c}^{(n)}$ vertices for all possible $n$. For example, in the case of i.i.d. initialized networks, the dependence of $g^{(4)}(x_{1},...,x_{4})$ on $G_{c}^{(4)}$ results in $g^{(4)}$ scaling as $1/N$ at leading order, while other connected correlation functions contribute to subleading effects at $O(1/N^{2})$ or less. More details on the Feynman rule prescription for general $g^{(r)}$, including the cases of non-i.i.d. initializations, can be found in Section 3.2 and 3.3 of Ref. [32].
@@ 4.1 Bayesian RG for NNGPs
At initialization, we take our network parameters to be normally distributed as [^20] We have introduced notation in which $i_{\ell}=1,...,n_{\ell}$ is an index for the space $O_{\ell}$ in a chosen basis. Thus, $b_{\ell}^{i_{\ell}}$ can be interpreted as the $i_{\ell}^{\rm th}$ bias parameter in the $\ell^{\rm th}$ layer.
@@
Here we have used the notation $y\sim\mathcal{N}(\mu,\sigma^{2})$ to mean that $y$ is a random variable sampled from a normal distribution with mean $\mu$ and variance $\sigma^{2}$. In words, (61) means that the parameters in each layer are each sampled from the same distribution (identically distributed), and that each layer is independent from every other layer. The final output of the network is given by
@@ 4.2.1 Implementation and network architecture
$\phi^{*}(x_{1},x_{2}):=(a-x_{1})^{2}+b(x_{2}-x_{1}^{2}),$ (91)
The input in this case is two-dimensional, $x\in I=\mathbb{R}^{2}$, and is sampled from the Gaussian distribution $X\sim\mathcal{N}(0,\mathbb{I}_{2})$, where $\mathbb{I}_{2}$ is the standard $2\times 2$ identity matrix. Training is performed such that the network output $\phi_{\theta}(x)$, where $\phi_{\theta}$ is the network function parameterized by $\theta$, minimizes the $L^{2}$ loss function,
@@
for the desired target, $y\in\mathbb{R}$, corresponding to the input, $x$. Optimization is performed using ‘vanilla’ gradient descent [^26] By ‘vanilla’ gradient descent, we mean traditional gradient descent as opposed to stochastic gradient descent..
It is typical in the training of NNs to start with a collection of random draws from a seed distribution over parameters, performing the chosen optimization procedure for each parameter draw in parallel. For our purposes, as was discussed in Section 2, it is natural to interpret the seed distribution as a prior over networks, and to treat the ensemble of parallel networks as realizations from the Bayesian posterior at each stage of training. In this way, we can approximate the posterior in a numerically tractable form by parsing calculable statistics from the ensemble. For our sample problem, we take the prior distribution to be Gaussian. Following the notation of Section 2, we denote the ensemble of parameters by $\{\theta_{n}(D)\}_{n=1}^{k}$, where $D=\{x_{i},\phi^{*}_{i}\}_{i\in\mathcal{I}}$ is a training sample. In this case, the training is supervised, so each network input datum, $x_{i}$, is matched with the corresponding ground truth model realization, $\phi_{i}\equiv\phi^{*}(x_{i})$.